### v2

@@ 3 Overview of techniques
Consider a utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, prices $\mathbf{p},\mathbf{q}\in\operatorname{pdom}(u)$ such that $\mathbf{q}\leq\mathbf{p}$. Then, $u(y)\geq\frac{1}{2}u(x)$ for any $y\in\mathcal{GD}^{u}(\mathbf{q},b)$ and $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$.
The second lemma shows that for every allocation of goods in the demand correspondence, there is a smaller or equal allocation of goods in the Gale demand correspondence. Although the first lemma was true for any utilities, this one critically relies on assuming the $\Sigma$-Gale-substitutes property.
@@
Let $u(\cdot)$ be a $\Sigma$-Gale-substitutes utility and let $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$ for $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ and $b>0$. Then, $\mathbf{p}\in\operatorname{pdom}(u)$, and there exists $y\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $y\leq x$.
To derive Theorem 1.3, we start with a competitive equilibrium $(\mathbf{x},\mathbf{p})$, and show that the dual of the Eisenberg–Gale program (5), $\varphi(\mathbf{q})$, has an optimal solution $\mathbf{q}^{\star}$ such that $\mathbf{q}^{\star}\leq\mathbf{p}$. The statement then follows from Lemma 3.1.
An intuitive argument would be the following: we start from $\mathbf{q}=\mathbf{p}$ where the total Gale demand is less than the supply (by Lemma 3.2), and use a tâtonnement-like procedure using the Gale-substitutes property. As long as $\mathbf{q}$ is not a minimizer of $\varphi(.)$, there will be an underdemanded item $j$ with respect to Gale demands. Reducing $p_{j}$ for such a good $j$ may only decrease the Gale demand for all agents on all other goods. This iterative process converges to an equilibrium point and, since the prices are only decreasing, we can conclude that the optimal limit price vector satisfies $\mathbf{q}^{\star}\leq\mathbf{p}$.
[…]
In Section 7, we prove Theorem 1.4, i.e., every separable utility function is $\Sigma$-Gale-substitutes. We show the theorem for general, not necessarily differentiable utility functions using subgradient arguments. We next outline the argument under the simplying assumption that utilities are differentiable and strictly concave. Let $u(y)=\sum_{j\in{\mathcal{M}}}v_{j}(y_{j})$, and let $\nabla_{j}u(y)=\partial v_{j}(y_{j})/\partial y_{j}$ denote the $j$-th component of the gradient. Then, the KKT conditions yield $\nabla_{j}u(y)\leq q_{j}u(y)/b$, with equality whenever $y_{j}>0$ (see Lemma 7.2). Let $y\in\mathcal{GD}^{u}(\mathbf{q},b)$ and $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ for $\mathbf{q}^{\prime}\leq\mathbf{q}$. Using the optimality conditions, it is easy to derive that $u(y^{\prime})\geq u(y)$. Consequently, for every good $j$ with $q_{j}=q^{\prime}_{j}$ and $y_{j}>0$, $\nabla_{j}u(y^{\prime})\geq\nabla_{j}u(y)$ follows. By strict concavity, we must have $y^{\prime}_{j}\leq y_{j}$ as required. The proof without the differentiability and strict concavity assumptions poses additional difficulties: for example, we would not be able to conclude $y^{\prime}_{j}\leq y_{j}$ at this point, and might need a specific choice of $y^{\prime}_{j}$ from $\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$.
Theorem 1.5 shows that generalized network utilities are Gale substitutes. To verify the Gale substitutes property, we need to take two price vectors $\mathbf{q}\geq\mathbf{q}^{\prime}$ and $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, and construct a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q^{\prime}},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q^{\prime}_{j}=q_{j}$; we have a second, analogous construction for budget changes. The proof is a finite algorithm that transforms the optimal generalized flow for demands $y$ to $y^{\prime}$ such that we throughout maintain primal-dual slackness condition corresponding to the Gale demand correspondence. The algorithm builds on the literature on maximum generalized flow algorithms. In particular, we use ideas from the classical highest gain augmenting path algorithm by Onaga Onaga (1966). While this does not give a polynomial-time algorithm, it was the starting point of further such algorithms, including Goldberg et al. (1991), Olver and Végh (2020), Végh (2014).
4 Efficiency of competitive equilibria
We now show the ‘price of anarchy’ result Theorem 1.1.
Theorem 4.1 (Restatement of Theorem 1.1).
Given a Fisher market instance, consider any competitive equilibrium $(\mathbf{x},\mathbf{p})$ and a Nash welfare maximizing allocation $\mathbf{y}$. Then,
$\left(\prod_{i\in{\mathcal{A}}}u_{i}(x_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}\geq\left(\frac{1}{\mathrm{e}}\right)^{\frac{1}{\mathrm{e}}}\left(\prod_{i\in{\mathcal{A}}}u_{i}(y_{i})^{b_{i}}\right)^{\frac{1}{\sum_{i\in{\mathcal{A}}}b_{i}}}.$
@@ 4 Efficiency of competitive equilibria
Without loss of generality, we can assume $\sum_{i\in{\mathcal{A}}}b_{i}=1$. Note that since $(\mathbf{x},\mathbf{p})$ is a competitive equilibrium, then $u_{i}(x_{i})\geq\min\left\{1,\frac{b_{i}}{\left\langle\mathbf{p},y_{i}\right\rangle}\right\}u_{i}(y_{i})$, as agent $i$ can use their budget, $b_{i}$, to buy at least $\frac{b_{i}}{\left\langle\mathbf{p},y_{i}\right\rangle}$ fraction of $y_{i}$. The lower bound holds because if $b_{i}<\left\langle\mathbf{p},y_{i}\right\rangle$, then $u_{i}\left(\frac{b_{i}}{\left\langle\mathbf{p},y_{i}\right\rangle}y_{i}\right)\geq\frac{b_{i}}{\left\langle\mathbf{p},y_{i}\right\rangle}u_{i}(y_{i})$ using concavity and $u_{i}(\mathbf{0})=0$. Thus,
@@
$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(y_{i})}{u_{i}(x_{i})}\leq\sum_{i\in{\mathcal{A}}}{b_{i}}\log\max\left\{1,\frac{\left\langle\mathbf{p},y_{i}\right\rangle}{b_{i}}\right\}.$
Let $b^{\prime}_{i}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}{\left\langle\mathbf{p},y_{i}\right\rangle}$, and let ${\mathcal{A}}^{\prime}$ be the set of agents such that $b^{\prime}_{i}\geq b_{i}$, $B\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in\mathcal{A}^{\prime}}b_{i}$, and $B^{\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in\mathcal{A}^{\prime}}b^{\prime}_{i}$. Therefore, using the log sum inequality,
$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(y_{i})}{u_{i}(x_{i})}\leq\sum_{i\in{\mathcal{A}}^{\prime}}{b_{i}}\log\frac{b^{\prime}_{i}}{b_{i}}\leq B\log\frac{B^{\prime}}{B}\,.$
By definition, $B\leq\sum_{i\in{\mathcal{A}}}b_{i}=1$. We also note that $B^{\prime}\leq\sum_{i\in{\mathcal{A}}}b_{i}^{\prime}=\sum_{i\in{\mathcal{A}}}\left\langle\mathbf{p},y_{i}\right\rangle\leq\sum_{j\in{\mathcal{M}}}p_{j}\leq\sum_{i\in{\mathcal{A}}}b_{i}=1$.
$\sum_{i\in{\mathcal{A}}}{b_{i}}\log\frac{u_{i}(y_{i})}{u_{i}(x_{i})}\leq\max_{B\leq 1}B\log\frac{1}{B}\leq\frac{1}{\mathrm{e}}\,.$
The statement follows. ∎
5 Properties of the Gale demand correspondence
In this section, we show some properties of the Gale demand correspondence, and also derive Theorem 1.2.
@@ 5 Properties of the Gale demand correspondence
Given any utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, for any $y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},y\right\rangle>b$, there exists $y^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},y^{\prime}\right\rangle\leq b$ and $b\log u(y^{\prime})-\left\langle\mathbf{q},y^{\prime}\right\rangle>b\log u(y)-\left\langle\mathbf{q},y\right\rangle$. In particular, if $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, then, $\left\langle\mathbf{q},y\right\rangle\leq b$.
@@
Let $y^{\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}y\cdot\frac{b}{\left\langle\mathbf{q},y\right\rangle}$. By concavity and $u(\mathbf{0})=0$, we get $u(y^{\prime})\geq\frac{b}{\left\langle\mathbf{q},y\right\rangle}u(y)$. Thus,
$b\log u(y^{\prime})-\left\langle\mathbf{q},y^{\prime}\right\rangle$ $\geq b\log u(y)-b\log\frac{\left\langle\mathbf{q},y\right\rangle}{b}-b>b\log u(y)-\left\langle\mathbf{q},y\right\rangle\,,$
Next, we derive some technical properties of the price domain $\operatorname{pdom}(u)$.

### v3

@@ 4 Nash welfare approximates competitive equilibrium
Given any utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, for any $\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},\mathbf{y}\right\rangle>b$, there exists $\mathbf{y}^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},\mathbf{y}^{\prime}\right\rangle\leq b$ and $b\log u(\mathbf{y}^{\prime})-\left\langle\mathbf{q},\mathbf{y}^{\prime}\right\rangle>b\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle$. In particular, if $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, then, $\left\langle\mathbf{q},\mathbf{y}\right\rangle\leq b$.
Proof.
Let $\mathbf{y}^{\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\mathbf{y}\cdot\frac{b}{\left\langle\mathbf{q},\mathbf{y}\right\rangle}$. By concavity and $u(\mathbf{0})=0$, we get $u(\mathbf{y}^{\prime})\geq\frac{b}{\left\langle\mathbf{q},\mathbf{y}\right\rangle}u(\mathbf{y})$. Thus,
$b\log u(\mathbf{y}^{\prime})-\left\langle\mathbf{q},\mathbf{y}^{\prime}\right\rangle$ $\geq b\log u(\mathbf{y})-b\log\frac{\left\langle\mathbf{q},\mathbf{y}\right\rangle}{b}-b>b\log u(\mathbf{y})-\left\langle\mathbf{q},\mathbf{y}\right\rangle\,$,
The next lemma follows directly from the subgradient Karush–Kuhn–Tucker (KKT) conditions used to characterize solutions in the Gale demand correspondence. For a detailed treatment of Lagrangian duality and KKT conditions for nondifferentiable functions, we refer the reader to Chapter 3 of Ruszczyński (2011).
@@
For a utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, let $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$. Then, there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(\mathbf{y})$ such that
$g_{j}\leq{q_{j}}\cdot\frac{u(\mathbf{y})}{b}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }y_{j}>0\,$,
Conversely, if such a $\mathbf{g}\in\partial^{\wedge}u(\mathbf{y})$ exists then $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$.
See 1.2
@@
By Lemma 4.1, all agents satisfy the budget constraints, $\left\langle\mathbf{q},\mathbf{y}_{i}\right\rangle\leq b_{i}$ for all $i$. Additionally, the market clears by the KKT conditions for (EG). For the approximate competitive equilibrium, we only need to show $u_{i}(\mathbf{y}_{i})\geq\frac{1}{2}\max_{\left\langle\mathbf{q},\mathbf{x}_{i}\right\rangle\leq b_{i}}u_{i}(\mathbf{x}_{i})$ for all agents.
Note that $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ is the allocation that maximizes Nash welfare, where each $\mathbf{y}_{i}$ maximizes $b_{i}\log u_{i}(\mathbf{y}_{i})-\left\langle\mathbf{q},\mathbf{y}_{i}\right\rangle$. By Lemma 4.2, there is a supergradient $\mathbf{g}_{i}$ such that $g_{ij}\leq q_{j}u_{i}(\mathbf{y}_{i})/b_{i}$ with equality whenever $y_{ij}>0$. Then, for any $\mathbf{x}_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},\mathbf{x}_{i}\right\rangle\leq b_{i}$,
$u_{i}(\mathbf{x}_{i})$ $\leq u_{i}(\mathbf{y}_{i})+\left\langle\mathbf{g}_{i},\mathbf{x}_{i}-\mathbf{y}_{i}\right\rangle$ (convexity of $u_{i}$)
@@
$\leq u_{i}(\mathbf{y}_{i})+\frac{u_{i}(\mathbf{y}_{i})}{b_{i}}\left\langle\mathbf{q},\mathbf{x}_{i}-\mathbf{y}_{i}\right\rangle$ ($g_{ij}\leq q_{j}u_{i}(\mathbf{y}_{i})/b_{i}$ with equality whenever $y_{ij}>0$)
$\leq\left(2-\frac{\left\langle\mathbf{q},\mathbf{y}_{i}\right\rangle}{b_{i}}\right)u_{i}(\mathbf{y}_{i})$ ($\left\langle\mathbf{q},\mathbf{x}_{i}\right\rangle\leq b_{i}$)
$\leq 2u_{i}(\mathbf{y}_{i})\,.$
This proves that $((\mathbf{y}_{i})_{i\in{\mathcal{A}}},\mathbf{q})$ forms a $2$-approximate competitive equilibrium. Additionally, the $2$-envy-freeness of the allocation follows directly from Lemma 2.10. ∎
One may wonder if the $2$-demand-approximate CE can be improved. The following example demonstrates that this bound of $2$ is tight.
Example 4.3 (Tight example on demand-approximate CE).
Consider the case with two agents and a single good. Both agents have unit budgets. The first agent’s utility for this good is linear:
$u_{1}(x_{1})=x_{1}.$
The second agent’s utility is defined as:
$u_{2}(x_{2})=\begin{cases}\alpha x_{2}&\text{for }x_{2}\leq\frac{1}{\alpha},\\ 1+\beta\left(x_{2}-\frac{1}{\alpha}\right)&\text{for }x_{2}>\frac{1}{\alpha},\end{cases}$
where $\alpha$ is a large number and $\beta\leq\alpha$.
When $\beta=\frac{1-\frac{1}{\alpha}}{1-\frac{2}{\alpha}}$, with the optimal dual variable (price) $q$ equal to $q=\frac{1}{x_{1}}=\frac{1}{1-\frac{1}{\alpha}}$, the Nash welfare maximizing allocation is
$x_{1}=1-\frac{1}{\alpha},\quad x_{2}=\frac{1}{\alpha}.$
Under this price, agent 2 can afford $x_{2}^{\prime}=1-\frac{1}{\alpha}$ with the budget, resulting in a utility of
$u_{2}(x_{2}^{\prime})=1+\beta\left(x_{2}^{\prime}-\frac{1}{\alpha}\right)=2-\frac{1}{\alpha}.$
In contrast, in the Nash welfare maximizing allocation, agent 2 only receives a utility of $1$.
Although Theorem 1.1 guarantees that the Nash welfare of any competitive equilibrium is at least a constant fraction of the optimum, the following example illustrates that agents can receive significantly different utilities across different equilibria.
Example 4.4 (Utility discrepancy across competitive equilibria).
Consider a market with $n$ agents and $n$ goods, where each agent has a unit budget. For $1\leq i\leq n-1$, agent $i$’s utility is given by
$u_{i}(\mathbf{x}_{i})=\min\{x_{ii},1\}+\varepsilon x_{in},$
for a small $\varepsilon>0$, and agent $n$’s utility is
$u_{n}(\mathbf{x}_{n})=x_{nn}.$
There exist at least two distinct competitive equilibria:
  • In the first equilibrium, all goods are priced at $1$, and each good $i$ is allocated entirely to agent $i$ for $1\leq i\leq n$. Agent $n$ thus attains a utility of $1$. This equilibrium is also a Nash welfare maximizing allocation.
  • In the second equilibrium, the first n1n-1 goods are priced at 0, while the $n$-th good is priced at $n$. Goods $1$ to n1n-1 are allocated to agents $1$ through n1n-1 respectively, and the $n$-th good is shared equally among all $n$ agents. As a result, agent $n$ obtains utility equal to $\frac{1}{n}$.
5 Agent-wise utility guarantee
We now turn to the proof of the main result of this paper, Theorem 1.3, asserting that for Gale-substitutes utility functions, each agent’s utility in the Nash welfare maximizing allocation is at least half of the maximum utility they could achieve in any competitive equilibrium. The proof relies on two key lemmas derived using Lagrangian duality and supergradient arguments. The first is a simple observation, showing that the Gale demand solution for lower prices yields at least half the utility of the utility maximizing bundle at the current prices.
@@ 5 Agent-wise utility guarantee
Consider a utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, prices $\mathbf{p},\mathbf{q}\in\operatorname{pdom}(u)$ such that $\mathbf{q}\leq\mathbf{p}$. Then, for any $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$ and $\mathbf{x}\in{\mathcal{D}}^{u}(\mathbf{p},b)$, it holds that $u(\mathbf{y})\geq\frac{1}{2}u(\mathbf{x})$.
@@
By Lemma 4.2, there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(\mathbf{y})$ such that $g_{j}\leq q_{j}u(\mathbf{y})/b$ with equality whenever $y_{j}>0$. Therefore,
$u(\mathbf{x})$ $\leq u(\mathbf{y})+\left\langle\mathbf{g},\mathbf{x}-\mathbf{y}\right\rangle\leq u(\mathbf{y})+\frac{u(\mathbf{y})}{b}\left\langle\mathbf{q},\mathbf{x}-\mathbf{y}\right\rangle$
$\leq u(\mathbf{y})\left(1+\frac{\left\langle\mathbf{q},\mathbf{x}\right\rangle}{b}\right)\leq u(\mathbf{y})\left(1+\frac{\left\langle\mathbf{p},\mathbf{x}\right\rangle}{b}\right)\leq 2u(\mathbf{y})\,.$
The second lemma establishes that for every allocation of goods in the standard demand correspondence, there exists a smaller or equal allocation of goods in the Gale demand correspondence. While Lemma 5.1 holds for any utility functions, this result critically depends on the assumption that the utility function satisfies the $\Sigma$-Gale-substitutes property. The proof is given in Appendix B, transforming the KKT-conditions of the demand correspondence ${\mathcal{D}}^{u}(\mathbf{p},b)$ to the KKT-conditions of $\mathcal{GD}^{u}(\mathbf{p},b)$ (Lemma 4.2), by making use of property ($\Sigma$).